LASER SPECTROSCOPY OF 3ν, 5ν and 6ν VIBRATION - ROTATION BANDS OF HBr

\underline{N. NISHIMIYA}, T. YUKIYA AND M. SUZUKI, Dept. of Electronic Engineering, Tokyo Institute of Polytechnics, Iiyama 1583, Atsugi City, 243-02 Kanagawa, Japan.

The vibration - rotation spectrum of H^{79/81}Br have been studied using a GaAs semiconductor laser and a titanium sapphire ring laser in the near infrared region. The rotational fine structures up to J=15 could be detected using a 20 m absorption cell. About 100 lines have been assigned in v = 3 ← 0, 5 ← 0 and 6 ← 0 bands with a resolving power of 0.001 cm^{−1}. Spectroscopic constants have been calculated with a least-squares fitting procedure by using the spectroscopic constants of v = 0 reported by Lonardo and Fusina (1). The values in cm^{−1} are determined as below;

G_{3} − G_{0} = 7405.2618(8)  B_{3} = 7.656220(15)  D_{3} × 10^{4} = 3.3701(5)
7404.1937(10)  7.653951(18)  3.3673(5)
G_{5} − G_{0} = 11887.9321(5)  B_{5} = 7.190004(18)  D_{5} × 10^{4} = 3.3661(10)
11886.2940(5)  7.187967(18)  3.3640(10)
G_{6} − G_{0} = 13990.8191(10)  B_{6} = 6.953022(45)  D_{6} × 10^{4} = 3.3777(35)
13988.9406(3)  6.951116(15)  3.3782(13).

The numbers given in the upper rows are those of I^{79}Br. The Dunham’s coefficients of HBr have also been determined by a joint analysis with the constants reported by the references (1, 2 ).

(1 ) G. Di Lonardo, and L. Fusina, J. Mol. Spectrosc., 148, 86–92 (1991). (2 ) V. Braun, and P. F. Bernath, J. Mol. Spectrosc., 167, 282–287 (1994).